# FineFilter: A Fine-grained Noise Filtering Mechanism for Retrieval-Augmented Large Language Models

### v1

@@ Abstract
Retrieved documents containing noise will hinder Retrieval-Augmented Generation (RAG) from detecting answer clues, necessitating noise filtering mechanisms to enhance accuracy.Existing methods use re-ranking or summarization to identify the most relevant sentences, but directly and accurately locating answer clues from these large-scale and complex documents remains challenging. Unlike these document-level operations, we treat noise filtering as a sentence-level MinMax optimization problem: first identifying the potential clues from multiple documents using contextual information, then ranking them by relevance, and finally retaining the least clues through truncation. In this paper, we propose FineFilter, a novel fine-grained noise filtering mechanism for RAG consisting of a clue extractor, a re-ranker, and a truncator. We optimize each module to tackle complex reasoning challenges: (1) Clue extractor firstly uses sentences containing the answer and similar ones as fine-tuned targets, aiming at extracting sufficient potential clues; (2) Re-ranker is trained to prioritize effective clues based on the real feedback from generation module, with clues capable of generating correct answer as positive samples and others as negative; (3) Truncator takes the minimum clues needed to answer the question (truncation point) as fine-tuned targets, and performs truncation on the re-ranked clues to achieve fine-grained noise filtering. Experiments on three QA datasets demonstrate that FineFilter significantly outperforms baselines in terms of performance and inference cost. Further analysis on each module shows the effectiveness of our optimizations for complex reasoning  [^1] Our code is available at https://anonymous.4open.science/r/FineFilter-5BE0.
@@
Qianchi Zhang1,2, Hainan Zhang1thanks: Corresponding author., Liang Pang3, Hongwei Zheng2, Yongxin Tong1, Zhiming Zheng1, 1Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing, School of Artificial Intelligence, Beihang University, China, 2Beijing Academy of Blockchain and Edge Computing, China, 3Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China, {zhangqianchi, zhanghainan}@buaa.edu.cn
@@ 1 Introduction
Figure 1: An illustration of the challenge in locating accurate answer clues from retrieved documents. The baseline RECOMP and RichRAG pick an incorrect clue from the $1^{th}$ document, whereas our FineFilter relies on the extraction, reranking, and truncation to identify the correct clue from the $4^{th}$ document.
@@
Retrieval-Augmented Generation (RAG) has demonstrated impressive performance across various knowledge-intensive NLP tasks Chen et al. (2022); Huang et al. (2023); Gao et al. (2023), but its effectiveness heavily depends on the relevance of the retrieved documents Liu et al. (2024). When retrieved documents contain noise or irrelevant information Zhu et al. (2024), the generation model struggles to detect answer clues because noise interferes with self-attention’s ability to reason over the correct context. Therefore, it is crucial to filter out irrelevant and low-value contexts.
@@
Current noise filtering methods primarily utilize re-ranking Wang et al. (2025); Ke et al. (2024) or summarization Xu et al. (2024); Zhu et al. (2024) models to identify the most relevant sentences, aiming at increasing the information density for RAG reasoning. The former re-ranks retrieval results based on metrics such as answer contribution or user preference. The latter retains the query-relevant sentences through summarization models. However, directly and accurately locating answer clues from the retrieved documents remains challenging, especially in complex reasoning scenarios. As shown in Figure 1, all the five documents recalled from a retriever contain query-relevant information. Both baseline RichRAG and RECOMP select the relevant sentences from the $1^{th}$ document, yet they produce incorrect answers. This is because these documents contain a multitude of seemingly relevant but unhelpful noisy information. Such document-level filtering is too coarse and struggles to capture effective answer clues precisely. Therefore, a fine-grained operation is required to retain sufficient and effective context for RAG.
@@
We treat the fine-grained noise filtering as a sentence-level MinMax optimization problem. First, we leverage contextual information to identify potential answer clues, which form the maximal subset capable of answering the question. Then, we carefully compare and re-rank these clues based on their completeness and relevance to the query in order to move effective clues to the forefront. Finally, we retain only the most essential clues through truncation, with the goal of minimizing reasoning costs for RAG. As shown in Figure 1, our approach first identifies the potential clues with a red background, then re-ranks these clues, ultimately placing the correct answer clue at the top. Notably, the last three clues are redundant and should be filtered out to improve the information density of the reasoning clues for RAG.
@@
In this paper, we propose a novel fine-grained noise filtering mechanism for RAG, named FineFilter, consisting of a clue extractor, a re-ranker, and a truncator. It leverages the clue extractor and re-ranker to provide sufficient and effective reasoning clues to the generation model while employing the truncator to filter noise to reduce reasoning costs. We design three optimization strategies for each module to tackle complex reasoning challenges: (1) Clue extractor uses all sentences containing the answer and their similar sentences based on KNN clustering as fine-tuning targets, since we find that RAG requires more relevant contextual information to reason the correct answer for multi-hop questions. Thus, the fine-tuned extractor can extract sufficient potential clues for complex reasoning. (2) Re-ranker is trained to prioritize effective clues based on the real feedback from the generation module, with clues capable of generating correct answers as positive samples while others as negative. (3) Truncator takes the minimal number of clues (truncation point) required for RAG to generate correct answers as the fine-tuning target. Based on the predicted point, the re-ranked clues are truncated to achieve fine-grained noise filtering.
@@
We conduct experiments on three open-domain question answering datasets, i.e., NQ, TriviaQA, and HotpotQA. The experimental results show that, whether based on LLaMA3 or Mistral, FineFilter outperforms the baseline models in terms of performance while significantly reducing the context required for inference. Further analysis of each module demonstrates the effectiveness of our optimization strategies for complex reasoning.
@@
  • We frame noise filtering as a sentence-level MinMax optimization, where the extractor and re-ranker gather sufficient and effective reasoning clues, while the truncator filters out noise to reduce reasoning costs.
  • Three strategies tackle complex reasoning: KNN-based extractor gathers sufficient relevant context, while re-ranker and truncator adapt quickly and effectively to RAG systems using generator feedback.
  • Experiments on three datasets show that filtering out unimportant noisy sentences enhances inference performance and efficiency.
Figure 2: The architecture of FineFilter includes three modules: clue extractor, re-ranker, and truncator. The top displays their training strategies and annotated data, while the bottom shows the noise filtering during inference.
@@ 2 Related Work
Retrievers often fetch noisy content, reducing output accuracy, while overly long contexts further hinder model efficiency. To address these challenges, some researchers utilize re-ranking methods to prioritize more relevant sentences. RichRAG Wang et al. (2025) uses a generative list-wise ranker to generate and rank candidate documents, ensuring the answer is comprehensive and aligns with the model’s preferences. Ke et al. (2024) proposes a novel bridge mechanism to optimize the connection between retrievers and LLMs in retrieval-augmented generation, improving performance in question-answering and personalized generation tasks. However, reranking sentences may disrupt the original logical structure of the document and generate unfaithful clues.
Other researchers utilize abstractive or extractive summarization models to identify query-relevant answer clues. Xu et al. (2024) propose leveraging LLMs as abstractive filters to compress retrieved text by targeting the most relevant sentences. Zhu et al. (2024) apply the information bottleneck principle to filter noise, striving to strike a balance between conciseness and correctness. Despite its potential benefits, this method is associated with high computational complexity during the training process, posing additional challenges for practical implementation. Xu et al. (2024); Wang et al. (2023) explore extractive filters to select the most relevant sentences. While these methods help eliminate irrelevant information, they also face the risk of over-compression, which may lead to a reduction in output accuracy. In another approach, Li et al. (2023) introduce the concept of Selective Context, which eliminates redundant content based on self-information metrics to enhance the efficiency of LLM inference. However, this technique may compromise the semantic coherence of the context.
@@ 3 Problem Formulation
where $\mathcal{S^{{}^{\prime}}}$ is the subset that is most capable of producing the correct answer, and $|\mathcal{S^{{}^{\prime}}}|$ is the number of sentences in $\mathcal{S^{{}^{\prime}}}$. The selection of $\mathcal{S}^{*}$ should dynamically adapt to the real feedback of a RAG system to balance informativeness and conciseness, ensuring a trade-off between computational efficiency and answer accuracy. The problem can be formalized as an NP-hard combinatorial optimization problem Wu et al. (2023), selecting the smallest, most relevant answer clues from a large set of documents to improve answer accuracy.
@@ 4 Methodology
In this section, we propose a three-stage noise filtering mechanism for RAG, called FineFilter, as shown in Figure 2. FineFilter consists of three modules: the clue extractor, the clue re-ranker, and the adaptive truncator. First, the clue extractor maximizes information gain to select potential answer clues from multiple documents, reducing the search space and improving the relevance of the candidate set. Next, the clue re-ranker optimizes the ranking of sentences using pairwise loss, ensuring the most relevant clues are prioritized. Finally, the adaptive clue truncator truncates the minimal necessary context, ensuring a balance between computational efficiency and answer accuracy.
Figure 3: The Exact Match performance of LLaMA3-8B-Instruct on three QA datasets between the top-5 documents and answer-containing sentences. The answer-containing sentences refer to all sentences in the same top-5 documents where the ground-truth answer appears, regardless of their query relevance.
@@ 4.1 Clue Extractor
The goal of clue extractor is to identify potential answer clues from multiple documents and construct a smaller query-relevant candidate set to reduce search space. We compare the performance of answer-containing sentences and the original retrieved documents in downstream tasks, as shown in Figure 3. We can see that filtering out the low-value information from the documents benefits RAG reasoning. Although not all answer-containing sentences are query-relevant, they approximate the maximal subset capable of addressing user queries and can serve as the optimization target for the clue extractor.
To optimize the sentence extraction process, we first introduce the concept of information gain. Given a query $q$ and a set of candidate sentences $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$, the information gain $\text{IG}(q,s_{i})$ of sentence $s_{i}$ is defined as:
@@
where $H(q)$ represents the entropy of the query $q$, measuring the uncertainty of the query; and $H(q\mid s_{i})$ represents the uncertainty of the query given the sentence $s_{i}$. In question answering tasks, information gain measures the reduction of uncertainty in the query by including a particular sentence. Typically, sentences that contain the answer directly reduce the unresolved part of the query, helping the model better understand the core of the query and improve the accuracy of the downstream generation module.
Based on this information gain concept, we first extract sentences from the retrieved document collection that contain the ground-truth answer as extraction targets. Given the query $q$, the ground-truth answer $y$ and retrieved sentences $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$, the answer-containing sentences is defined as:
@@
Then, we finetune an LLM model as the clues $Extractor$ to generate answer-containing sentences $\mathcal{S}^{a}$ based on the query $q$ and the retrieved sentences $\mathcal{S}$ with a specific prompt(see Appendix A.1). The loss function of $Extractor$ model is defined as:
@@
Finally, our clue extractor has the ability to generate the potential candidate clues based on the user query and retrieved documents in the inference:
@@
We find that answer-containing sentences significantly improve performance on the simple QA dataset, i.e., NQ, but less so on the complex QA dataset, i.e., TriviaQA and HotpotQA, as shown in Figure 3. Therefore, we propose a KNN-based similar sentence extraction strategy for complex reasoning scenarios. For simple questions, we directly select sentences containing the answer as the extractor’s optimization targets, as these sentences provide the key information to answer the question and significantly reduce the uncertainty. For more complex questions, we first select sentences containing the answer and then further select sentences semantically similar to the answer using the K-Nearest Neighbors (KNN) method Guo et al. (2003). Although these sentences may not directly contain the answer, they provide contextual information related to the question’s answer, helping the model better understand the nature of the question and generate a more accurate answer. We utilize both the answer-containing sentences and the KNN-based similar sentences as the extractor’s optimization targets. This KNN-based strategy allows the system to respond to queries of varying complexity flexibly, ensuring higher accuracy while minimizing computational overhead.
@@ 4.2 Clue Reranker
The sentences selected by the clue extractor often contain multiple relevant clues, but their relevance may vary, requiring further re-ranking. To achieve this, we train a re-ranker using pairwise loss to optimize the ranking model, ensuring that the most relevant sentences are ranked at the front.
@@
We use the real RAG-generated feedback to annotate the training data for the re-ranker, as the QA performance on complex questions heavily depends on the characteristics of the generation module. First, we pair each of the extracted clue sentences $s_{j}^{c}\in\mathcal{S}^{c}$ with the query $q$ as $(s_{j}^{c},q)$, where sentence $s_{j}^{c}$ that enables the downstream generation module to produce the correct answer for $q$ is considered as positive sample $s_{\text{positive}}$, while other sentences are treated as negative samples $s_{\text{negative}}$. [^2] If no candidate clues can generate the correct answer, or if all samples can generate the correct answer, the sample will be removed from the annotated data. The goal of $Reranker$ is to minimize the following pairwise loss function Karpukhin et al. (2020) to improve the relevance ranking:
@@
where $\text{sim}(q,*)$ is the semantic similarity between the query $q$ and the sentence * by $Reranker$ model. By minimizing this loss function, the $Reranker$ model can effectively identify the most relevant clues and prioritize them accordingly.
@@
Given the query $q$ and the extracted sentences $\mathcal{S}^{c}$, $Reranker$ model calculate the relevance score between every sentence $s_{j}^{c}\in\mathcal{S}^{c}$ and query $q$. The re-ranked answer clues are defined as:

### v4

@@ Abstract
Retrieved documents containing noise will hinder Retrieval-Augmented Generation (RAG) from detecting answer clues, necessitating noise filtering mechanisms to enhance accuracy. Existing methods use reranking or summarization to identify the most relevant sentences, but directly and accurately locating answer clues from these large-scale and complex documents remains challenging. Unlike these document-level operations, we treat noise filtering as a sentence-level MinMax optimization problem: first identifying potential clues from multiple documents, then ranking them by relevance, and finally retaining the minimum number of clues through truncation. In this paper, we propose FineFilter, a novel fine-grained noise filtering mechanism for RAG, consisting of a clue extractor, a reranker, and a truncator. We optimize each module to tackle complex reasoning challenges: (1) The clue extractor first uses sentences containing the answer and similar ones as fine-tuning targets, aiming to extract sufficient potential clues; (2) The reranker is trained to prioritize effective clues based on the real feedback from the generation module, with clues capable of generating correct answers as positive samples and others as negative; (3) The truncator takes the minimum number of clues needed to answer the question (truncation point) as fine-tuning targets, and performs truncation on the reranked clues to achieve fine-grained noise filtering. Experiments on three QA datasets demonstrate that FineFilter significantly improves QA performance over baselines on both LLaMA3 and Mistral. Further analysis confirms its effectiveness in complex reasoning, robustness to unreliable retrieval, and generalization to different scenarios.
@@
Qianchi Zhang1, Hainan Zhang1thanks: Corresponding author., Liang Pang3, Ziwei Wang1, Hongwei Zheng2, Yongxin Tong1, Zhiming Zheng1, 1Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing, School of Artificial Intelligence, Beihang University, China, 2Beijing Academy of Blockchain and Edge Computing, China, 3Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China, {zhangqianchi, zhanghainan}@buaa.edu.cn
@@ 1 Introduction
Figure 1: An illustration of the challenge in locating accurate answer clues from retrieved documents. Both baselines RECOMP and RichRAG select an incorrect clue from the $1^{st}$ document, whereas our FineFilter identifies the correct clue from the $4^{th}$ document via extraction, reranking, and truncation.
@@
Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023) has demonstrated impressive performance across various NLP tasks Chen et al. (2022); Huang et al. (2023); Gao et al. (2023), but its effectiveness heavily depends on the relevance of retrieved documents Yu et al. (2023); Zhang et al. (2024a). When retrieved documents contain noise or irrelevant information Shi et al. (2023); Liu et al. (2024), the generation model struggles to detect answer clues because noise interferes with self-attention’s ability to reason over the correct context. Therefore, it is crucial to filter out irrelevant and low-value contexts.
@@
Current noise filtering methods primarily utilize reranking Wang et al. (2025); Ke et al. (2024); Qin et al. (2024) or summarization Xu et al. (2024); Zhu et al. (2024) models to identify the most relevant sentences, aiming at increasing the information density for RAG reasoning. The former reranks the retrieval results based on metrics such as answer contribution or user preference Zhu et al. (2023). The latter retains the query-relevant sentences through summarization models. However, directly and accurately locating answer clues from the retrieved documents remains challenging, especially in complex reasoning scenarios. As shown in Figure 1, all the five documents retrieved by the retriever contain query-relevant information. Both baselines RichRAG Wang et al. (2025) and RECOMP Xu et al. (2024) select the relevant sentences from the $1^{st}$ document, yet fail to generate correct answers. This is because these documents contain a multitude of seemingly relevant but unhelpful noisy information. Such document-level filtering is too coarse and struggles to capture effective answer clues precisely. Therefore, a fine-grained operation is required to retain sufficient and effective context for RAG.
@@
We treat fine-grained noise filtering as a sentence-level MinMax optimization problem. First, we leverage contextual information to identify potential answer clues, which form the maximal subset capable of answering the question. Then, we carefully compare and rerank these clues based on their completeness and relevance to the query in order to move effective clues to the forefront. Finally, we retain only the most essential clues through truncation, with the goal of minimizing the necessary contexts for RAG. As shown in Figure 1, our approach first identifies the potential clues with a red background, then reranks these clues, ultimately placing the correct answer clue at the top. Notably, the last three clues are redundant and should be filtered out to improve the information density of the reasoning clues for RAG.
@@
In this paper, we propose a novel fine-grained noise filtering mechanism for RAG, named FineFilter, consisting of a clue extractor, a reranker, and a truncator. It leverages the clue extractor and the reranker to provide sufficient and effective reasoning clues to the generation model while employing the truncator to filter noise to reduce reasoning contexts. We design three optimization strategies for each module to tackle complex reasoning challenges: (1) The clue extractor uses all sentences containing the answer and their similar sentences based on K-Nearest Neighbors (KNN) clustering Guo et al. (2003); Peterson (2009) as fine-tuning targets, since we find that RAG requires more relevant contextual information to reason the correct answer for multi-hop questions. Thus, the fine-tuned extractor can extract sufficient potential clues for complex reasoning. (2) The reranker is trained to prioritize effective clues based on the real feedback from the generation module, with clues capable of generating correct answers as positive samples and others as negative. (3) The truncator takes the minimal number of clues (truncation point) required for RAG to generate correct answers as the fine-tuning target. Based on the predicted point, the reranked clues are truncated to achieve fine-grained noise filtering.
@@
Experimental results on three open-domain QA datasets (NQ, TriviaQA, and HotpotQA) show that FineFilter, whether based on LLaMA3 or Mistral, consistently outperforms baseline models in terms of performance while significantly reducing the context required for inference. Further analysis confirms its effectiveness in complex reasoning, robustness to unreliable retrieval, and generalization to different scenarios.
@@
  • We frame noise filtering as a sentence-level MinMax optimization, where the extractor and reranker gather sufficient and effective reasoning clues, while the truncator filters out noise to improve reasoning density.
  • Three strategies tackle complex reasoning: KNN-based extractor gathers sufficient relevant context, while reranker and truncator adapt quickly and effectively to RAG systems using generator feedback.
  • Experiments on three datasets show that filtering out unimportant noisy sentences enhances inference performance.
Figure 2: The architecture of FineFilter consists of three modules: clue extractor, reranker, and truncator. The top displays their training strategies and annotated data, while the bottom shows the noise filtering during inference.
@@ 2 Related Work
Retrievers often fetch noisy content, reducing output accuracy, while overly long contexts further hinder model efficiency Xia et al. (2024). To address these challenges, some researchers utilize reranking methods to prioritize more relevant sentences. RichRAG Wang et al. (2025) uses a generative list-wise ranker to generate and rank candidate documents, ensuring the answer is comprehensive and aligns with the model’s preferences. Ke et al. (2024) propose a novel bridge mechanism to optimize the connection between retrievers and LLMs in retrieval-augmented generation, improving performance in question-answering and personalized generation tasks. CPC Liskavets et al. (2025) ranks sentence relevance with context-aware embeddings. However, reranking sentences may disrupt the original logical structure of the document and generate unfaithful clues.
Other researchers utilize abstractive or extractive summarization models to identify query-relevant answer clues. Xu et al. (2024) propose leveraging LLMs as abstractive filters to compress retrieved text by targeting the most relevant sentences. Zhu et al. (2024) use the information bottleneck principle to balance conciseness and correctness, but this incurs significant training complexity. Xu et al. (2024); Wang et al. (2023) propose extractive filters to select relevant sentences, which reduce irrelevant information but may cause over-compression and harm accuracy. Li et al. (2023) enhance inference efficiency by removing redundant content based on self-information, though this may affect semantic coherence. AdaComp Zhang et al. (2024b) adaptively adjusts document-level compression but applies relatively mild compression overall.
@@ 3 Problem Formulation
where $\mathcal{S^{{}^{\prime}}}$ is the subset that is most capable of producing the correct answer, and $|\mathcal{S^{{}^{\prime}}}|$ is the number of sentences in $\mathcal{S^{{}^{\prime}}}$. The selection of $\mathcal{S}^{*}$ should dynamically adapt to the real feedback of a RAG system to balance informativeness and conciseness. The problem can be formalized as an NP-hard combinatorial optimization problem Wu et al. (2023), selecting the smallest, most relevant answer clues from a large set of documents to improve answer accuracy.
@@ 4 Methodology
In this section, we introduce FineFilter, a three-stage noise filtering mechanism for RAG, as illustrated in Figure 2. FineFilter comprises three modules: a clue extractor, a clue reranker, and an adaptive truncator. The clue extractor selects potential answer clues from multiple documents by maximizing information gain, reducing the search space and enhancing candidate relevance. The reranker then applies pairwise loss to prioritize the most relevant clues. Finally, the adaptive truncator selects the minimal necessary context to increase inference density, thus improving answer accuracy.
Figure 3: The Exact Match performance of LLaMA3-8B-Instruct on three QA datasets, comparing Top-5 retrieved documents with all sentences containing the ground-truth answer from those documents, regardless of query relevance.
@@ 4.1 Clue Extractor
The clue extractor aims to identify potential answer clues from multiple documents to construct a smaller, query-relevant candidate set, thereby reducing the search space. As shown in Figure 3, we compare downstream performance using answer-containing sentences versus full retrieved documents. Filtering out low-value content enhances RAG reasoning. While not all answer-containing sentences are query-relevant, they approximate the maximal subset capable of answering queries and thus serve as a suitable optimization target.
To guide the extraction of informative sentences, we first introduce the concept of information gain. Given a query $q$ and a set of candidate sentences $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$, the information gain $\text{IG}(q,s_{i})$ of sentence $s_{i}$ is defined as:
@@
where $H(q)$ denotes the entropy of the query $q$, quantifying its inherent uncertainty; and $H(q\mid s_{i})$ represents the uncertainty of the query given the sentence $s_{i}$. In question answering tasks, information gain reflects how much a sentence reduces uncertainty about the query. Typically, sentences containing the answer directly reduce the unresolved part of the query, helping the model better focus on the core intent and improving the accuracy of the downstream generation module.
Based on the concept of information gain, we first extract sentences from the retrieved document collection that contain the ground-truth answer as targets for extraction. Given a query $q$, the ground-truth answer $y$ and retrieved sentences $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$, the answer-containing sentences are defined as:
@@
Then, we fine-tune an LLM as the clue $Extractor$ to generate answer-containing sentences $\mathcal{S}^{a}$ based on the query $q$ and the retrieved sentences $\mathcal{S}$ with a specific prompt (see Appendix B.2). The loss function of $Extractor$ model is defined as:
@@
Finally, the clue extractor is capable of generating the potential candidate clues based on the user query and retrieved documents during inference:
@@
We observe that answer-containing sentences significantly improve performance on simple QA datasets, i.e., NQ, but yield smaller gains on complex ones, i.e., TriviaQA and HotpotQA, as shown in Figure 3. Therefore, we propose a KNN-based strategy for extracting semantically similar sentences in complex reasoning scenarios. For simple questions, we directly select answer-containing sentences as the extractor’s optimization targets, as they provide essential information and help reduce uncertainty. For more complex questions, we first select sentences containing the answer and then further select sentences semantically similar to these answer-containing sentences using the KNN method. Although these sentences may not contain the answer directly, they offer contextual clues that help the model better interpret the question and produce a more accurate response. We use both answer-containing and KNN-based similar sentences as the extractor’s optimization targets, allowing the extractor to adapt to queries of varying complexity while improving accuracy and reducing necessary contexts.
@@ 4.2 Clue Reranker
The clue extractor often selects sentences with multiple relevant clues of varying importance, necessitating reranking. We address this by training a reranker with pairwise loss to prioritize the most relevant sentences.
@@
We use the real RAG-generated feedback to annotate the training data for the reranker, as the QA performance on complex questions heavily depends on the characteristics of the generation module. First, we pair each of the extracted clue sentences $s_{j}^{c}\in\mathcal{S}^{c}$ with the query $q$ as $(s_{j}^{c},q)$, where sentence $s_{j}^{c}$ that enables the downstream generation module to produce the correct answer for $q$ is considered as positive sample $s_{\text{positive}}$, while other sentences are treated as negative samples $s_{\text{negative}}$ [^1] If no candidate clues can generate the correct answer, or if all samples can generate the correct answer, the sample will be removed from the annotated data.. The $Reranker$ aims to minimize the following pairwise loss function Karpukhin et al. (2020) to improve relevance ranking:
@@
where $\text{sim}(q,*)$ represents the semantic similarity between the query $q$ and the sentence * by $Reranker$ model. Minimizing this loss function enables the $Reranker$ model to effectively identify and prioritize the most relevant clues.
@@
Given the query $q$ and the extracted sentences $\mathcal{S}^{c}$, the $Reranker$ model calculates the relevance score between every sentence $s_{j}^{c}\in\mathcal{S}^{c}$ and query $q$. The reranked clues are defined as: